Subject: 3.5-m Users Committee minutes 10/06/08
From: Bruce Gillespie
Submitted: Wed, 8 Oct 2008 13:49:11 -0600
Message number: 1115
(previous: 1114,
next: 1116
up: Index)
APO 3.5-m Users Committee Phonecon, 10/06/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, Al
Harper, Meredith Drosback (for Remy Indebetouw), Mark Klaene, Russet
McMillan, Guy Stringfellow (for John Bally), Michael Strauss
Absent: Jon Holtzman, Jon Fulbright
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
The only substantive comment on user feedback came from Suzanne
Hawley. She and her students have noticed that in TUI, the windowing
function for DIS applies the same way for both the blue and red
images--they can't be set separately, and the spectra appear at
different places on the blue and red chips. At a minimum, this
"feature" should be highlighted in the users documentation, and
possibly we should consider revising the software so that the DIS blue
and red images can be windowed separately.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark pointed to the recent spate of good
weather and a well-behaved telescope. Now that the future of SDSS-III
is secured, we are beginning to spend funds that had been held back on
needed site improvements. We have continued diagnostic and repair
work on the tertiary rotation system, and its geared encoder is now
working better. The tertiary homing problem is on the schedule for
repair during October. Current performance anomalies with DIS cryo
and vacuum systems are being watched, and are hinting at future
maintenance. TripleSpec experienced a vacuum gauge failure, which is
being replaced and an ion pump is being added to the instrument;
TripleSpec should be ready for scheduled observations next week. For
CIF projects, we are close to permanently moving Agile to the TR2
port, with its new rotator. The new axes controllers are getting
close to being put on the telescope, and planning for next year's
direct-drive motor replacements in alt and az is underway.
Guy asked about the status of the loose wire recently found hanging
into SPIcam's field of view. Mark said that the loose wire was being
re-secured. Suzanne added a comment that as a part of the move of
Agile to a new mirror-cell port, Agile instrument control will be
implemented in TUI, probably by the end of the year.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 9/6/08
through 10/1/08
Mark Klaene
0) Overview
We have finally seen the end to the summer monsoons--the humidity has
dropped and clearer skies are now the norm, and fall is definitely in
the air. The telescope has been working quite well, taking good
advantage of the clear nights. On site-related matters, we are
nearing the completion of the installation of a new boiler for Dorm B
and repair of the fire alarm system replacement will begin this week.
1) Telescope
Most of the problems with tertiary rotation have been fixed with the
differential wiring modification done during the September engineering
time. New power supplies are on order to address the axial stepper
motor noise, and these will be installed during the October
engineering time.
2) Instruments
The DIS blue ion pump continues to run at high current, and we may
need to open the dewar in the near future for servicing. DIS red and
SPIcam are fully operational with no problems. Service was required
on the DIS red cryotiger, which was accomplished during a recent
period of poor weather.
Echelle and NIC-FPS remained cold and powered on during the shutdown.
We are waiting for UVa to design and build a new NIC-FPS fan-out board
that will hopefully reduce readout noise.
TripleSpec had to be warmed up due to a vacuum gauge failure, with the
repair still in progress. Most of the TripleSpec spare parts are now
at the site, and final instrument acceptance is pending the delivery
of documentation.
Agile has a new mask to reduce the light scattered around the CCD.
The instrument rotator for the TR2 port is now in testing in the lab.
GFP is currently warm since it is not on the schedule until late
November. APOLLO remains operational, and has had a recent record-
breaking run of of lunar-return photons.
3) CIF projects
New software was created for the altitude and az axis controllers, and
is being tested on the rotation axis. Planning for next years CIF
projects is in work. The major CIF project for next year is direct
drives for azimuth and altitude axes.
**********************************
Calibration (Flat, Arc lamps) upgrades - Klaene:
Mark mentioned that we have been looking at the suite of flat-field
and emission-line calibration lamps on the telescope in the context of
meeting users requirements for our new and upgraded instrumentation.
Bill Ketzeback has been researching this topic, and we have identified
possible improvements aimed at increasing the flux in the far blue/
near UV for DIS flat fields, and in providing additional lines for
wavelength calibration in this part of the spectrum. We are
interested to hear if the other optical instruments, and especially
the IR instruments, have adequate flat field and wavelength
calibration capability from our present lamps or if improvements are
desired. Other improvements are possible and may be worth
considering, such as more lamps, projector systems, fiber-fed cal
sources, internal cal sources, etc. Please contact us if you would
like more information on the current system and possible upgrades.
But in order to proceed, we need user feedback; Suzanne asked for an
action item.
ACTION: The Users Committee representatives are to ask their
respective users for suggestions on improving the calibration lamp
system on the telescope, e.g., flux, wavelength distribution, emission-
line calibration, etc., and report at the next meeting.
**********************************
CIF plans for next year - Hawley, Klaene:
The major CIF project for next year is the direct-drive replacement
for the alt and az telescope axes. There may be time and resources
for other smaller projects (e.g., cal lamps, as above) and minor
upgrades. Suzanne encourages the Users Committee members to talk with
their users about telescope and instrument upgrades. There is some
"opportunity space" here, particularly for instrumentation upgrades.
**********************************
Science highlights due - Hawley:
For those institutions that have not submitted their annual science
highlights to Suzanne and Bruce, the window for acceptance is rapidly
closing.
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: Committee members should send one or two science highlights
from the telescope users from their institutions to Suzanne (slh at
astro.washington.edu) and Bruce (gillespi at apo.nmsu.edu) not later
than 1 October for use in the Annual Report to the ARC Board of
Governors meeting in November.
STATUS: A few institutions have not yet turned in their highlights.
[new actions from this month]:
ACTION: The Users Committee representatives are to ask their
respective users for suggestions on improving the calibration lamp
system on the telescope, e.g., flux, wavelength distribution, emission-
line calibration, etc., and report at the next meeting.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 3 November, at
8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1115 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO